OpenStudy (bellezabrilla):

Write a polynomial equation for each situation. A bedroom has a length of x+3 feet and a width of x feet. Find the area when x=10. Please help !!!!

OpenStudy (bellezabrilla):

by plugging in 10 which would be 10+3

OpenStudy (bellezabrilla):

@3mar

OpenStudy (mathmale):

the bedroom floor forms a rectangle. Draw this rectangle and label the length and width, please. Use the Draw utility, below.

zepdrix (zepdrix):

You do you remember how to find the area of a rectangle? :)

OpenStudy (bellezabrilla):

yea \[a= l w\]

OpenStudy (bellezabrilla):

@zepdrix

zepdrix (zepdrix):

Ok good. And they told us our length and width,\[\large\rm \color{orangered}{\ell=x+3}\qquad\qquad\qquad \color{royalblue}{w=x}\]So we'll plug this information in for our L and W in our area formula,\[\large\rm A=\color{orangered}{\ell}\color{royalblue}{w}\]\[\large\rm A=\color{orangered}{(x+3)}\color{royalblue}{x}\]Make sense?

OpenStudy (bellezabrilla):

okay thank you so much i have a better understanding now

zepdrix (zepdrix):

`Write a polynomial equation for each situation.` They want you to get rid of the brackets and write it as a polynomial before you plug x=10 in. Do you understand how to distribute the x to each term in the brackets?

OpenStudy (bellezabrilla):

no

zepdrix (zepdrix):

Multiply the x to each term in the brackets. \(\large\rm A=(x+3)\color{royalblue}{x}=x\color{royalblue}{x}+3\color{royalblue}{x}\) Another way we can write xx is using exponent notation. \(\large\rm A=x^2+3x\)